how many Br atoms are in 2.50g of Br
1
Expert's answer
2019-05-16T06:24:56-0400

Bromine exists as a molecule Br2. Calculating first number of molecules Br2 in 2.50g of bromine:


"N(Br_2) = \frac{m(Br_2)}{M(Br_2)}\cdot N_A = \frac{2.50g}{160\frac{g}{mol}}\cdot 6.02\times 10^{23}mol^{-1} = 9.406\times 10^{21}"


Number of Br atoms is two times more than number of Br2 molecules:


"N(Br) = 2N(Br_2) = 2 \cdot 9.406 \times 10^{21} = 1.88 \times 10^{22}"